SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (Nov.8, 2005)

Attendees: Rainer Bartoldus, Prafulla Behera, Kris Chaisanguanthum, Al Eisner, Fang Fang, Selina Li, Rahmat Rahmat, Dong Su, Dan Walker.

Operations

  • BaBar has been taking cosmics with magnetic field since Monday swing shift and expects to see beams on Sunday.

EMT

  • Dan is working on OPR monitoring for EMT related to prescaling. It was concluded that we don't need to monitoring presacle in JAS, only in fastmon for TRG experts. Rainer suggested to test L3 event stream (offline) for prescaling.
  • We are not exactly sure if it was FEX or GLINK that causes deadtime for EMT.

DCT/DCZ

  • The new versions of the TSF firmware didn't work for production boards, but were okay with the prototype boards. The memory used is different for prototype and production boards. Other than that, it's not clear why there is a difference between the production and the prototype versions. The DAQ firmware still gets the header, but no TSF segments seen. We don't see segments sent to ZPD neither. It was suspected that something broke in the engine. Rainer suggested Marc to re-route the old version of the firmware to make sure the routing software still works fine.

GLT

  • A timing shift of clock32 is seen in DRC. This can be from master crate (we are not in PEP clock yet). The finest resolution we can see in TRG is clock15, so it cannot be due to Trigger.
  • Single photon group would like to have a meeting on Friday to present their physics cuts proposal. Al suggested that this would fit better into a Trigger/Filter AWG meeting.

Online

  • In case of the unphysical SL and cell ever occur again, the segments are now flagged with 255 highest bit and will still get shipped out for further diagnosis. Rainer suggested to ship input container and flag the event to be looked at later. We will try to avoid printing any statements from the ROM.

L3

  • We are still planning on merging the L3 meeting to the TRG operation meeting, given that all L3 people show up to the operation meeting.
  • Rainer has always wanted to have the trigger mask controled by trgconfig rather than run control. This will be implemented soon.
  • Every once a while, L3PCASERVER would pick up the wrong trgconf and so failed after calibrations were taken. There wasn't a simple solution, however, there is a chance that it might get resolved by having a faster farm machine to handle processing the calibration data so that the server won't get stuck with the wrong configuration.
  • Prafulla will try to get some time to work on the OPR plots this week.
  • After this TRG operation meeting, L3 and fastmon folks stayed and discussed the situation with L3 JAS plots. We went through the plots in JAS now, and found some plots that can be removed from there. We decided to clean out the directory in L3TMonTools and to stop using the script to automatically generate ~275 xml files.

Data Quality and Monitoring

  • Somehow Jim Hamilton had an older version of fastmon running and created some file access problem for fastmon .ps archive this morning. It has been fixed.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.